# huff: A Python package for Market Area Analysis

### v1

@@
Market area models are used in economic geography, regional economics, spatial planning, geoinformation science, and marketing, enabling the analysis and forecasting of market areas and customer flows for retail and service locations. The classical and most popular approach is the Huff model (Huff 1962, 1963, 1961) and its numerous derivates and extensions, such as the Multiplicative Competitive Interaction (MCI) Model (Nakanishi and Cooper 1974, 1982). Typical research applications include examining the influence of store attributes and transport costs on consumer store choice, forecasting the revenue of new locations, or predicting the impact of new locations on existing ones (De Beule et al. 2014; Fittkau 2004; Li and Liu 2012; Mensing 2018; Oruc and Tihi 2012; Suárez-Vega et al. 2015; Wieland 2015, 2019).
In health geography, such models are used to analyse catchment areas with respect to medical practices and hospitals (Bai et al. 2023; Fülop et al. 2011; Jia 2016; Latruwe et al. 2023; Rhein et al. 2025; Wieland 2018), and they are also increasingly being linked to methods for analyzing the supply structure and accessibility of health locations (Liu L 2022; Rauch et al. 2023; Subal J 2021). Moreover, market area models are also applied to other location-related contexts such as airports or recreation facilities (Wang et al. 2022; Wang et al. 2026).
@@
  • The calibration of the Huff model based on observed data on consumer behavior and/or store sales is difficult because the model is nonlinear in its weighting parameters (Huff 2003; Wieland 2017). In this context, the MCI model (Nakanishi and Cooper 1974, 1982) has been developed as an econometric estimation technique based on a linearization (log-centering transformation). As this approach requires empirical market shares for fitting, it is applied in cases where customer-store interaction data was obtained by surveys or secondary data (Baviera-Puig et al. 2016; Latruwe et al. 2023; Oruc and Tihi 2012; Suárez-Vega et al. 2015; Wieland 2015, 2019). Several other researchers developed and used nonlinear iterative fitting approaches, especially when no empirical customer-store interactions are available, but only total sales of the locations investigated (De Beule et al. 2014; Güßefeldt 2002; GH et al. 1972; Li and Liu 2012; Liang et al. 2020; Mensing 2018; Orpana and Lampinen 2003; Wieland 2017). Due to the pronounced sensitivity of market area models to weighting schemes, the availability of multiple calibration approaches is essential in market area analysis.
@@
Due to data confidentiality requirements, the early development of the huff library took place in a private repository; the public repository was initialized more recently to provide open access for reproducibility and review. The huff Python package has been publicly developed and published via the [Python Package Index] (https://pypi.org/project/huff/) since April 2025. As of submission, it has undergone 41 releases, showing continuous improvement and feature additions. The library is actively used: since its first release (version 1.0.0) in April 2025, it has been downloaded approximately 21,900 times from the Python Package Index (source: pepy.tech, accessed February 14, 2026).

### v2

@@
Market area models are used in economic geography, regional economics, spatial planning, geoinformation science, and marketing, enabling the analysis and forecasting of market areas and customer flows for retail and service locations. The classical and most popular approach is the Huff model (Huff 1962, 1963, 1964) and its numerous derivates and extensions, such as the Multiplicative Competitive Interaction (MCI) Model (Nakanishi and Cooper 1974, 1982; Cooper and Nakanishi 1983). See the “Mathematical appendix” section for the structure of the mentioned models. Typical research applications include examining the influence of store attributes and transport costs on consumer store choice, forecasting the revenue of new locations, or predicting the impact of new locations on existing ones (De Beule et al. 2014; Fittkau 2004; Li and Liu 2012; Mensing 2018; Oruc and Tihi 2012; Suárez-Vega et al. 2015; Wieland 2015, 2018a).
In health geography, such models are used to analyse catchment areas with respect to medical practices and hospitals (Bai et al. 2023; Fülop et al. 2011; Jia 2016; Latruwe et al. 2023; Rhein et al. 2025; Wieland 2018b), and they are also increasingly being linked to methods for analyzing the supply structure and accessibility of health locations (Liu et al. 2022; Rauch et al. 2023; Subal et al. 2021). Moreover, market area models are also applied to other location-related contexts such as airports or recreation facilities (Wang et al. 2022; Wang et al. 2026).
@@
  • The calibration of the Huff model based on observed data on consumer behavior and/or store sales is difficult because the model is nonlinear in its weighting parameters (Huff 2003; Wieland 2017). In this context, the MCI model (Nakanishi and Cooper 1974, 1982; Cooper and Nakanishi 1983) has been developed as an econometric estimation technique based on a linearization (log-centering transformation). As this approach requires empirical market shares for fitting, it is applied in cases where customer-store interaction data was obtained by surveys or secondary data (Baviera-Puig et al. 2016; Latruwe et al. 2023; Oruc and Tihi 2012; Suárez-Vega et al. 2015; Wieland 2015, 2018a). Several other researchers developed and used nonlinear iterative fitting approaches, especially when no empirical customer-store interactions are available, but only total sales of the locations investigated (De Beule et al. 2014; Güßefeldt 2002; Haines Jr et al. 1972; Li and Liu 2012; Liang et al. 2020; Mensing 2018; Orpana and Lampinen 2003; Wieland 2017). Due to the pronounced sensitivity of market area models to weighting schemes, the availability of multiple calibration approaches is essential in market area analysis.
@@
Due to data confidentiality requirements, the early development of the huff library took place in a private repository; the public repository was initialized more recently to provide open access for reproducibility and review. The huff Python package has been publicly developed and published via the [Python Package Index] (https://pypi.org/project/huff/) since April 2025. As of submission, it has undergone 44 releases, showing continuous improvement and feature additions. The library is actively used: since its first release (version 1.0.0) in April 2025, it has been downloaded approximately 23,700 times from the Python Package Index (source: pepy.tech, accessed March 02, 2026).
@@
Mathematical appendix
In the basic Huff model, the utility of supply location $j$ for the customers in customer origin $i$, $U_{ij}$ is (Huff 1962):
$U_{ij}=A_{j}^{\gamma}t_{ij}^{-\lambda}$
where $A_{j}$ is the attraction (size) of supply location $j$, $t_{ij}$ is the travel time from $i$ to $j$, and $\gamma$ and $\lambda$ are weighting parameters.
Given $I$ customer origins (i=1,2,,Ii=1,2,...,I) and $J$ supply locations (j=1,2,,Jj=1,2,...,J), the interaction probability (or market share) of origin $i$ with respect to location $j$, equals:
$p_{ij}=\frac{U_{ij}}{\sum_{j=1}^{J}U_{ij}}$
The expected customer or expenditure flows from customer origin $i$ to supply location $j$, $E_{ij}$, is:
$E_{ij}=p_{ij}C_{i}$
where $C_{i}$ is the customer or expenditure potential in origin $i$.
The total market area of location $j$, $T_{j}$, equals (Huff 1964):
$T_{j}=\sum_{i=1}^{I}E_{ij}$
The Multiplicative Competitive Interaction Model is formalized as follows (Nakanishi and Cooper 1974):
$p_{ij}=\frac{\prod_{h=1}^{H}A_{h_{j}}^{\gamma_{h}}}{\sum_{j=1}^{J}\prod_{h=1}^{H}A_{h_{j}}^{\gamma_{h}}}$
where $A_{h_{j}}$ is the $h$-th characteristic of supplier $j$, and $\gamma_{h}$ is the corresponding weighting coefficient.
The log-centering transformation of the MCI model equals (Nakanishi and Cooper 1974):
$\log\left(\frac{p_{ij}}{\widetilde{p}_{i}}\right)=\sum_{h=1}^{H}\hat{\gamma}_{h}\log\frac{A_{h_{j}}}{\widetilde{A}_{h_{j}}}+\log\left(\frac{\epsilon_{ij}}{\widetilde{\epsilon}_{i}}\right)$
where $\widetilde{p}_{i}$, $\widetilde{A}_{h_{j}}$, and $\widetilde{\epsilon}_{i}$ are the geometric means of $p_{ij}$, $A_{h_{j}}$, and $\epsilon_{ij}$, with $\epsilon_{ij}$ being the stochastic error term.
The inverse log-centering transformation is (Nakanishi and Cooper 1982):
$\hat{p}_{ij}=\frac{\exp{\sum_{h=1}^{H}\hat{\gamma}_{h}\log\frac{A_{h_{j}}}{\widetilde{A}_{h_{j}}}}}{\sum_{j=1}^{J}\exp{\sum_{h=1}^{H}\hat{\gamma}_{h}\log\frac{A_{h_{j}}}{\widetilde{A}_{h_{j}}}}}$
where $\hat{p}_{ij}$ is the estimated interaction probability (or market share) of origin $i$ with respect to location $j$.